OpenStudy (anonymous):

Multiply 3i(8-i). Subtract 4-(-2+3i)-(-4-i).

OpenStudy (anonymous):

for the first one what do you think?

OpenStudy (anonymous):

3+24i. But why?

OpenStudy (anonymous):

because 3i*8 = 24i and 3i*(-i)=-3i^2 and i^2=-1 so 3i*(-i)=3 and therefore you have 24i+3

OpenStudy (anonymous):

\[3\iota \left( 8-\iota \right) =24\iota-3\iota ^{2}=24\iota-3\left( -1 \right)=24\iota+3 \]

OpenStudy (anonymous):

for the second one distribute the - sign to the terms in te parentheses and then get rid of the parentheses and simply add and subtract

OpenStudy (anonymous):

10-2i

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Thank you both very much.